Solve the following equation with the method of your choice:

y/3y-4 = 2/y+4

Answer by stanbon(75887)

Solve the following equation with the method of your choice:
y/3y-4 = 2/y+4
----
Cross-multiply to get:
y(y+4) = 2(3y-4)
---
y^2+4y = 6y-8
---
y^2 -2y + 8 = 0
(y-4)(y+2) = 0
y = 4 or y = -2
=====================
Cheers,
Stan H.
=====================